coffee grounds inc has two limds of coffee. Coffee A costs $9 per kg and coffee B costs $6 per kg. how many kilograms of each should be combined to obtain 150 kg of a blend worth $8 per kg

Answer by mananth(16949)

cost ---------------- quantity
Coffee A 6 ----------------x lb
Coffee B 9 ----------------150-x lb
Total 8.00% ---------------- 150

6x+9(150-x)=8*150

6x+1350 -9x=1200
6x-9x=1200-1350
-3x=-150
/-3
x=50 lb $6.00
100 lb $9
m.ananth@hotmail.ca